OpenStudy (anonymous):

Need help with this 3x-5y=-8 y=x

OpenStudy (anonymous):

you have a system of equation where: Y = X We can use this to find either x or y on the first equation, which ever you find better: 3x - 5y = -8 (where y = x so,) 3y - 5y = -8 -2y = -8 y = (8/2) y = 4 if y = 4 and x = 4 we know that: y = 4 x = 4